OpenStudy (anonymous):

graph the line 5x+2y-8=0

hartnn (hartnn):

have you ever before graphed the line ? or found the intercepts ?

OpenStudy (anonymous):

no i need help changing it to the y=mx+b format. I get really confused with these problems.

hartnn (hartnn):

ok, after changing it to y=mx+b format, will u be able to graph the line ?

OpenStudy (anonymous):

yea

hartnn (hartnn):

to get that into y=mx+b format you just need to isolate y you can start by adding 8 on both sides, what u get ?

OpenStudy (anonymous):

Yea I get 5x+2y+8 right

OpenStudy (anonymous):

*=

hartnn (hartnn):

yes, 5x+2y=8 is correct now there's also x term with the y 5x is added to it, so next step would be to SUBTRACT 5x from both sides, what u get ?

OpenStudy (anonymous):

2y+3x?

OpenStudy (anonymous):

=

hartnn (hartnn):

nopes... you can't subtract a constant term with x term 5x+2y = 8 subtracting 5x from both sides, 2y = 8-5x got this ?

OpenStudy (anonymous):

Or 2y+8-5x?

OpenStudy (anonymous):

Oh okay and after that I should subtract 2 from each side?

OpenStudy (anonymous):

y=-5x+6?

hartnn (hartnn):

nopes, 2 is MULTIPLIED with y so to get rid of it, you should DIVIDE 2 on both sides... makes sense ?

OpenStudy (anonymous):

y=-5x+4?

hartnn (hartnn):

now divide 2 on both sides of 2y =8-5x

hartnn (hartnn):

divide 2 on each term

OpenStudy (anonymous):

y=-5/2x+4?

hartnn (hartnn):

yes! correct :)

hartnn (hartnn):

thats in y=mx+b format :)

OpenStudy (anonymous):

okay thanks can you help me with factoring?

hartnn (hartnn):

sure

OpenStudy (anonymous):

I understand some but not all. For example, p^2+7p+10

hartnn (hartnn):

find 2 numbers whose sum is +7 and product is 10

OpenStudy (anonymous):

5 and 2?

hartnn (hartnn):

correct! so, break up 7x as 5x+2x

hartnn (hartnn):

x^2 +5x+2x+10 = 0 now what can you factor out from 1st 2 terms ?

OpenStudy (anonymous):

Okay this is where I get very confused.

hartnn (hartnn):

ok, x^2 mean x*x right ?

OpenStudy (anonymous):

Yes

hartnn (hartnn):

so, what do u see common in x*x + 5*x ??

OpenStudy (anonymous):

They all have an x?

hartnn (hartnn):

yup, so factor out x you would do x^2 + 5x = x (x+5) got this ?

OpenStudy (anonymous):

Yes

hartnn (hartnn):

good, so now what about 2x+10 ?? note 10 = 2*5

OpenStudy (anonymous):

I would divide 2?

hartnn (hartnn):

when you see a common factor, just bring that out...like this 2x+2*5 2 is common so, 2x+10 = 2 (x+5) got it ?

OpenStudy (anonymous):

Yes so it would be (5+p)(2+p)?

hartnn (hartnn):

good! absolutely correct :)

OpenStudy (anonymous):

Okay Thanks! I have one more problem if it's not a bother?

hartnn (hartnn):

no bother.... just that, next time, ask new Q in new post :)

OpenStudy (anonymous):

Alright. And so the problem is a little long, it's an exponential problem. I usually get these but this one is different. The world population in 2005 was 6.2 billion and growing exponentially at a rate of 1.14% per year. the function P(t)=6.2(10^0.005t) provides a good model for the population growth pattern. In this model what does P(t) represent and what does t represent

OpenStudy (anonymous):

Also says Explain how you can be sure that this function givesthe correct popuation in 2005

hartnn (hartnn):

eh....i am not so good at this, can you ask this in new post...some1 will surely help u :)

OpenStudy (anonymous):

Alright thanks for everything though.